# Let’s Get to the Point: LLM-Supported Planning, Drafting, and Revising of Research-Paper Blog Posts

### v2

@@ 4.2.4. Study Procedure
5. Evaluation Studies’ Results
5.1. Editing Power
5.1.1. Lab Study - Exploratory
Participants made greater change to the initial LLM draft blog post when using Papers-to-Posts as compared to the baseline tool. For each condition, we calculated the Levenshtein distance between participants’ initial LLM draft blog posts and their blog posts at the point at which they would feel comfortable publicly publishing the blog post. Levenshtein distance measures how many character insertions, deletions, and replacements would be needed to transform one text into another. For statistical testing, we employed a paired-samples t-test and observed that this Levenshtein distance is significantly higher with Papers-to-Posts (M=3977.05, SD=1146.46) compared to the baseline (M=2764.15, SD=1042.16) (Paired-Samples t-Test, t(19)=4.31, p¡.0005). [^16] We note that this significance test was not part of our initial hypotheses and was therefore exploratory in nature. The difference in Levenshtein distance does not appear attributable to the time spent on the task or the change in blog post length. Participants did not spend substantially more time to reach the point at which they were comfortable publicly publishing the blog post using Papers-to-Posts as compared to the baseline (see Section 5.5). Figure 5a shows Levenshtein distance plotted against the time taken to get to a publishable blog post, with the removal of one outlier. Furthermore, participants changed the length of their blog posts, measured in characters, less with Papers-to-Posts (M=-162.00, Q1=-1126.25, Q3=297.00) than with the baseline (M=1971.00, Q1=965.25, Q3=2514.25) (Appendix 26(b)). [^17] As you can see in Appendix 26(b), the initial length of treatment drafts was significantly higher than the initial length of baseline drafts.
5.1.2. Deployment Study - RQ1
We investigated if the increase in editing power observed in the lab study would be reflected in a more realistic setting and if Papers-to-Posts required fewer writing actions (i.e., inserting/deleting a character or text span in the draft or LLM instructions) [^18] We did not count when participants moved, inserted, or deleted a blog post section in Papers-to-Posts as a writing action because these actions involved clicking a button rather than manipulating text. in order to achieve a given amount of change in the blog post. To analyze editing power, we excluded the three study runs in which the participant made no change to the initial draft. We also excluded six study runs in which the participant claimed to have made edits to the blog post outside the tool, beyond small edits like those related to grammar, formatting, and adding images. The resulting analysis thus included 28 study runs (treatment: 14; baseline: 14).
Participants using Papers-to-Posts took fewer writing actions to create one unit of change in Levenshtein distance to the blog post (Figure 5(c)). Papers-to-Posts also led to a greater final Levenshtein distance for a given amount of active time, which we define as time during which participants were taking any actions in the tool (Figure 5(b)). Once more, this is despite the fact that participants did not spend substantially more active time using Papers-to-Posts (M=25.71 minutes, SD=14.72) than the baseline (M=35.93 minutes, SD=34.38) (Appendix 27(a)). [^19] Considering that there were three outliers with a high amount of active time in the baseline condition, we can also look at the median behavior. We see little difference between the conditions (treatment: M=22.00, Q1=17.25, Q3=34.00; baseline: M=22.50, Q1=16.00,27.00). Furthermore, participants changed the length of their blog posts, measured in characters, less using Papers-to-Posts less using Papers-to-Posts (M=-5.50, Q1=-525.25, Q3=961.00) than using the baseline (M=710.00, Q1=47.25, Q3=4261.25) (Appendix 27(b)). [^20] As you can see in Appendix 27(b), the initial length of treatment drafts was significantly higher than the initial length of baseline drafts.
5.1.3. Summary
In the lab study, our exploratory analysis revealed that Papers-to-Posts led to increased editing power, in the sense that participants made more change to their blog posts in a given amount of time. In the deployment study, participants experienced the same outcome in a more realistic setting. We further found that participants made more change to their blog posts for a given number of writing actions.
@@ 5.2.1. Lab Study - H1
To compare responses under the two conditions in the within-subjects lab study, we employed a non-parametric test, given the data’s Likert nature. Furthermore, we used a sign test rather than a Wilcoxon signed rank test, which is normally used to analyze non-parametric within-subjects data for two conditions of one factor, due to the violation of that test’s assumption of a symmetric distribution of the paired differences about the median. Participants were significantly more satisfied with the final blog post generated with Papers-to-Posts (M=6.00, Q1=5.00, Q3=7.00) as compared to the baseline tool (M=5.00, Q1=5.00, Q3=6.00) (Dependent-Samples Sign Test, S=11, p¡.05). The majority of participants (11) were more satisfied in the treatment condition (equally: 7, less: 2).
Inspecting participants’ Likert responses regarding aspects of their blog post satisfaction, we see that participants particularly demonstrated more satisfaction with the content covered in the Papers-to-Posts blog post (M=6.00, Q1=5.00, Q3=7.00) compared to the baseline blog post (M=5.50, Q1=4.75, Q3=6.00) (Figure 6b). The majority of participants (14) were more satisfied with the content under the treatment condition (equally: 2, less: 4). Although it is not apparent from the medians, participants were overall more satisfied with the blog post structure using Papers-to-Posts (M=6.00, Q1=5.00, Q3=6.25) than the baseline (M=6.00, Q1=5.00, Q3=6.00) (Figure 6d). The plurality of participants (9) were more satisfied with the structure under the treatment (equally: 6, less: 5). Participants did not show increased satisfaction with the blog post’s tone under the treatment (M=6.00, Q1=5.00, Q3=6.00) compared to the baseline (M=5.50, Q1=5.00, Q3=6.00) (Figure 6c). A plurality of participants were not more satisfied with the tone under the treatment (more: 7, equally: 6, less: 7).
Looking at participants’ perceived failure to complete their task (Figure 7c), we observe that participants had lower perceived failure with Papers-to-Posts (M=2.00, Q1=2.00, Q3=3.00) than with the baseline (M=3.00, Q1=2.00, Q3=4.00). The plurality of participants (9) thought that they were more successful in writing a research-paper blog post with Papers-to-Posts compared to the baseline (equally: 7, less: 4). This further corroborates participants’ blog post satisfaction under the treatment condition.
@@ 5.2.2. Deployment Study - RQ2a
In the between-subjects deployment study, we observed no difference in the median satisfaction under Papers-to-Posts (M=6.00, Q1=5.00, Q3=6.00) versus the baseline (M=6.00, Q1=5.25, Q3=6.00) in blog post satisfaction (Figure 6j). Again with respect to median satisfaction, participants were equally satisfied with the blog post content (treatment: M=6.00, Q1=5.00, Q3=6.50; baseline: M=6.00, Q1=3.50, Q3=6.00; Figure 6k), tone (treatment: M=6.00, Q1=5.00, Q3=6.00; baseline: M=6.00, Q1=5.00, Q3=7.00; Figure 6l), and structure (treatment: M=6.00, Q1=6.00, Q3=6.00; baseline: M=6.00, Q1=5.25, Q3=7.00; Figure 6m) when using Papers-to-Posts versus the baseline. In terms of perceived failure to complete their task (Figure 7h), participants had equal median perceived failure with Papers-to-Posts (M=3.00, Q1=2.00, Q3=3.00) and the baseline (M=3.00, Q1=2.00, Q3=3.00).
@@ 5.2.3. Summary
We observed a significant increase in blog post satisfaction under the treatment condition during the lab study and found that participants were particularly more satisfied with the blog post content, suggesting that Papers-to-Posts best addressed its design goal to support planning blog post content (DG1). We did not observe the same increases in blog post satisfaction in the deployment study. This discrepancy may be related to the fact that participants were not writing under a time constraint in the deployment study. With sufficient time to write the blog post to their liking (only three blog posts were not shared with others), participants may not have noticed a pronounced difference in blog post quality between conditions.
@@ 5.3. Tool Satisfaction
For measuring tool satisfaction, we used the five-question Likert scale on tool acceptance from Kocielnik et al. (Kocielnik et al., 2019) (Figure 6e-i). The five responses were averaged to obtain one score.
@@ 5.3.1. Lab Study - H2
We used the non-parametric sign test to evaluate the lab study data, as it was Likert data and violated the Wilcoxon signed rank test’s assumption of a symmetric distribution of the paired differences about the median. While participants had a higher median satisfaction with Papers-to-Posts (M=5.90, Q1=5.20, Q3=6.50) than the baseline tool (M=5.60, Q1=4.80, Q3=6.00), this difference was not significant (Dependent-Samples Sign Test, S=13, p=n.s.).
@@ 5.3.2. Deployment Study - RQ2b
Similarly to the lab study, we observed slightly greater satisfaction with the tool under the treatment condition (M=5.60, Q1=5.00, Q3=6.40) than the baseline condition (M=5.40, Q1=4.05, Q3=5.80) (Figure 6n-r). Based on median results, participants under the treatment condition were more likely to find the tool helpful (treatment: M=6.00, Q1=5.00, Q3=6.50; baseline: M=5.00, Q1=5.00, Q3=6.00), use it if it were available (treatment: M=6.00, Q1=5.00, Q3=7.00; baseline: M=5.00, Q1=4.00, Q3=6.00), and recommend the tool (treatment: M=6.00, Q1=5.00, Q3=6.50; baseline: M=5.00, Q1=3.25, Q3=6.00). They were also less likely to find the tool annoying (treatment: M=2.00, Q1=2.00, Q3=4.00; baseline: M=2.50, Q1=2.00, Q3=4.00). On the other hand, they were equally likely under both conditions to find the tool useful for productivity (treatment: M=6.00, Q1=4.50, Q3=7.00; baseline: M=6.00, Q1=4.00, Q3=6.00).
@@ 5.3.3. Summary
Across both studies, we observed higher satisfaction with Papers-to-Posts than the baseline tool, though this difference was not significant according to the lab study results. In particular, participants found Papers-to-Posts more helpful.
@@ 5.4. Cognitive Load
To measure cognitive load, we used the NASA TLX Index questionnaire (Hart and Staveland, 1988) and summed the responses for the mental demand, temporal demand, effort, and frustration questions (Figure 7a-b,d-e) [^21] See 5.2 for discussion of the responses to the performance question.
@@ 5.4.1. Lab Study - H3
We utilized the non-parametric sign test to evaluate the lab study data, as it was Likert data and violated the Wilcoxon signed rank test’s assumption of a symmetric distribution of the paired differences about the median. We saw no significant difference between the baseline and treatment conditions (Dependent-Samples Sign Test, S=8, p=n.s.), although participants reported slightly lower cognitive load with Papers-to-Posts (M=13.00, Q1=11.50, Q3=16.30) than the baseline tool (M=14.50, Q1=12.00, Q3=18.30).
We saw the largest difference in median results with respect to perceived effort (treatment: M=3.50, Q1=3.00, Q3=5.00; baseline: M=5.00, Q1=3.75, Q3=5.00). The majority of participants (11) reported using less effort with Papers-to-Posts than with the baseline tool (more: 4, equal: 5). Meanwhile, Papers-to-Posts led to an increase in perceived mental demand (treatment: M=5.00, Q1=3.00, Q3=5.00; baseline: M=4.00, Q1=3.75, Q3=5.25). The small plurality of participants (9) perceived more mental demand with Papers-to-Posts (less: 8, equal: 3). We saw no change in the median result for perceived temporal demand (treatment: M=3.00, Q1=2.00, Q3=4.25; baseline: M=3.00, Q1=2.00, Q3=4.25) or frustration (treatment: M=2.00, Q1=1.00, Q3=4.00; baseline: M=2.00, Q1=2.00, Q3=3.00). An equal number of participants found the treatment more (9) versus less (9) temporally demanding, and the small plurality of participants (9) found the treatment less frustrating (more: 7, equal: 4).
@@ 5.4.2. Deployment Study - RQ2c
Once again, we found slightly lower cognitive load under the treatment condition (M=15.00, Q1=9.50, Q3=17.50) in comparison to the baseline condition (M=15.50, Q1=10.75, Q3=18.75) (Figure 7f-g,i-j). Looking at the breakdown of questions, the treatment led to less perceived mental demand (M=4.00, Q1=2.00, Q3=5.50) than the baseline (M=5.00, Q1=3.25, Q3=5.75) but higher perceived effort (treatment: M=5.00, Q1=3.00, Q3=6.00; baseline: M=4.50, Q1=3.00, Q3=5.00). We found no change in the median result for perceived temporal demand (treatment: M=3.00, Q1=2.00, Q3=4.00; baseline: M=3.00, Q1=2.00, Q3=4.00) or frustration (treatment: M=3.00, Q1=2.00, Q3=5.00; baseline: M=3.00, Q1=2.00, Q3=4.75).

### v3

@@ 5.3.3. Summary
There are three scatterplots labeled ”a,” ”b,” and ”c.” For each scatterplot, there is a legend titled ”Condition” with a blue dot labeled ”baseline” and an orange dot labeled ”treatment.” The ”a” scatterplot has an x-axis labeled ”Time (minutes)” with tick marks for 20 through 60 and a y-axis labeled ”Levenshtein Distance” with tick marks for 1000 through 8000. The plot shows an orange line going from around (20,3000) on the left to around (60,4250) on the right and a blue line going from around (22,2000) on the left to around (60,3000) on the right. Shading around each line overlaps more towards the start of the lines. The shading of one line overlaps with another line in only one area: the orange shading overlaps with the start of the blue line. The ”b” scatterplot has an x-axis labeled ”Active Time (minutes)” with tick marks for 0 through 120 and a y-axis labeled ”Levenshtein Distance” with tick marks for 0 through 12000. The plot shows an orange line going from around (5,2500) on the left to around (60,5500) on the right and a blue line going from around (10,1500) on the left to around (120,9000) on the right. Shading around each line overlaps more towards the start of the lines. The shading of one line slightly overlaps with another line in two areas: the orange shading overlaps with the start of the blue line and the blue shading overlaps with near the start of the orange line. The ”c” scatterplot has an x-axis labeled ”Writing Actions” with tick marks for 0 through 8000 and a y-axis labeled ”Levenshtein Distance” with tick marks for 0 through 10000. The plot shows an orange line going from around (0,3000) on the left to around (3000,6000) on the right and a blue line going from around (0,1500) on the left to around (10000,8500) on the right. Shading around each line overlaps towards the start of the lines. The shading of one line slightly overlaps with another line in only one area: the blue shading overlaps with near the start of the orange line.
@@
The top of the figure shows a legend with the following mapping: dark red = strongly disagree, medium red = disagree, light red = somewhat disagree, gray = neither agree nor disagree, light blue = somewhat agree, medium blue = agree, and dark blue = strongly agree. Below, there are two areas. The left area is labeled ”Lab Study,” and the right area is labeled ”Deployment Study.” In each area, there are nine questions labeled ”a” through ”i” for the lab study and ”j” through ”r” for the deployment study. The questions are: 1) ”I am satisfied with the overall quality of the blog post.”; 2) ”I am satisfied with the content covered in the blog post.”; 3) ”I am satisfied with the tone of the blog post.”; 4) ”I am satisfied with the structure of the blog post.”; 5) ”I found the tool to be helpful.”; 6) ”I would use the tool if it were available.”; 7) ”I would recommend the tool to my friends and colleagues.”; 8) ”I found the tool to be able to improve my productivity.”; 9) ”I found the tool to be annoying or distracting.” Below each question, there are two horizontal stacked bars, one labeled ”treatment” and the other ”baseline.” In the lab study area, treatment is shown to perform somewhat better than baseline for each question except the last, which is ambiguous. For each question, the medians for treatment followed by baseline are: 1) agree vs somewhat agree, 2) agree vs between somewhat agree and agree, 3) agree vs between somewhat agree and agree, 4) agree vs agree, 5) agree vs between somewhat agree and agree, 6) agree vs agree, 7) between somewhat agree and agree vs between somewhat agree and agree, 8) agree vs between somewhat agree and agree, and 9) disagree vs disagree. In the deployment study area, treatment is shown to perform somewhat better than baseline for the second question and fifth through ninth questions. For the first, third, and fourth question, it is ambiguous. For each question, the medians for treatment followed by baseline are: 1) agree vs agree, 2) agree vs agree, 3) agree vs agree, 4) agree vs agree, 5) agree vs somewhat agree, 6) agree vs somewhat agree, 7) agree vs somewhat agree, 8) agree vs agree, and 9) disagree vs between disagree and somewhat disagree.
@@
The top of the figure shows a legend with the following mapping: dark red = 1 (very low), medium red = 2, light red = 3, gray = 4, light blue = 5, medium blue = 7, and dark blue = 7 (very high). Below, there are two areas. The left area is labeled ”Lab Study,” and the right area is labeled ”Deployment Study.” In each area, there are five questions labeled ”a” through ”e” for the lab study and ”f” through ”j” for the deployment study. The questions are: 1) ”Mental Demand (How mentally demanding was the task?)”; 2) ”Temporal Demand (How hurried or rushed was the pace of the task?)”; 3) ”Performance (How unsuccessful were you in accomplishing what you were asked to do?)”; 4) ”Effort (How hard did you have to work to accomplish your level of performance?)”; and 5) ”Frustration (How insecure, discouraged, irritated, stressed, and annoyed were you?).” Below each question, there are two horizontal stacked bars, one labeled ”treatment” and the other ”baseline.” In the lab study area, treatment is shown to perform better than baseline for effort and somewhat better for performance, while baseline is shown to perform somewhat better than baseline for temporal demand and frustration. For mental demand, it is ambiguous which performed better. For each question, the medians for treatment followed by baseline are: 1) 5 vs 4, 2) 3 vs 3, 3) 2 vs 3, 4) between 3 and 4 vs 5, and 5) 2 vs 2. In the deployment study area, treatment is shown to perform somewhat better than baseline for mental demand. For the rest which did better is ambiguous. For each question, the medians for treatment followed by baseline are: 1) 4 vs 5, 2) 3 vs 3, 3) 3 vs 3, 4) 5 vs between 4 and 5, and 5) 3 vs 3.
@@ 5.4.1. Lab Study - H1
To compare responses under the two conditions, we employed a non-parametric test, given the data’s Likert nature. Furthermore, we used a sign test rather than a Wilcoxon signed rank test, which is normally used to analyze non-parametric within-subjects data for two conditions of one factor, due to the violation of that test’s assumption of a symmetric distribution of the paired differences about the median. Participants were significantly more satisfied with the final blog post generated with Papers-to-Posts (M=6.00, Q1=5.00, Q3=7.00) as compared to the baseline tool (M=5.00, Q1=5.00, Q3=6.00) (Dependent-Samples Sign Test, S=11, p¡.05). The majority of participants (11) were more satisfied in the treatment condition (equally: 7, less: 2).
Inspecting participants’ Likert responses regarding aspects of their blog post satisfaction, we see that participants particularly demonstrated more satisfaction with the content covered in the Papers-to-Posts blog post (M=6.00, Q1=5.00, Q3=7.00) compared to the baseline blog post (M=5.50, Q1=4.75, Q3=6.00) (Figure 7b). The majority of participants (14) were more satisfied with the content under the treatment condition (equally: 2, less: 4). The median participant was only slightly more satisfied with the blog post’s tone under the treatment (M=6.00, Q1=5.00, Q3=6.00) compared to the baseline (M=5.50, Q1=5.00, Q3=6.00) (Figure 7c). With respect to the blog post’s structure, the median participant was equally satisfied using Papers-to-Posts (M=6.00, Q1=5.00, Q3=6.25) compared to the baseline (M=6.00, Q1=5.00, Q3=6.00) (Figure 7d).
In addition, looking at participants’ perceived failure to complete their task (Figure 8(b)c), we observe that the median participant had lower perceived failure with Papers-to-Posts (M=2.00, Q1=2.00, Q3=3.00) than with the baseline (M=3.00, Q1=2.00, Q3=4.00). The plurality of participants (9) thought that they were more successful in writing a research-paper blog post with Papers-to-Posts compared to the baseline (equally: 7, less: 4). This further corroborates participants’ blog post satisfaction under the treatment condition.
@@ 5.4.2. Deployment Study - RQ2a
In the deployment study, we observed no difference in the median blog post satisfaction under Papers-to-Posts (M=6.00, Q1=5.00, Q3=6.00) versus the baseline (M=6.00, Q1=5.25, Q3=6.00) (Figure 7j). Again with respect to median satisfaction, participants were equally satisfied with the blog post content (treatment: M=6.00, Q1=5.00, Q3=6.50; baseline: M=6.00, Q1=3.50, Q3=6.00; Figure 7k), tone (treatment: M=6.00, Q1=5.00, Q3=6.00; baseline: M=6.00, Q1=5.00, Q3=7.00; Figure 7l), and structure (treatment: M=6.00, Q1=6.00, Q3=6.00; baseline: M=6.00, Q1=5.25, Q3=7.00; Figure 7m) when using Papers-to-Posts versus the baseline. In terms of perceived failure to complete their task (Figure 8(b)h), participants had equal median perceived failure with Papers-to-Posts (M=3.00, Q1=2.00, Q3=3.00) and the baseline (M=3.00, Q1=2.00, Q3=3.00).
@@ 5.4.3. Summary
We observed a significant increase in blog post satisfaction under the treatment condition during the lab study and found that participants were particularly more satisfied with the blog post content, suggesting that Papers-to-Posts best addressed its design goals to support understanding (DG1) and control (DG2) of the content included in the blog post. We did not observe the same increases in blog post satisfaction in the deployment study. This discrepancy may be related to the fact that participants were not writing under the pressure of a time constraint. With sufficient time to write the blog post to their liking (only three blog posts were not shared with others), participants may not have noticed a pronounced difference in blog post quality between conditions.
@@ 5.5. Tool Satisfaction
For measuring tool satisfaction, we used the five-question Likert scale on tool acceptance from Kocielnik et al. (Kocielnik et al., 2019) (Figure 7e-i). The five answers were averaged to obtain one score.
@@ 5.5.1. Lab Study - H2
We used the sign test to evaluate the lab study data, as it was non-parametric data and violated the Wilcoxon signed rank test’s assumption of a symmetric distribution of the paired differences about the median. While the median participant had a higher satisfaction with Papers-to-Posts (M=5.90, Q1=5.20, Q3=6.50) than the baseline tool (M=5.60, Q1=4.80, Q3=6.00), this difference was not significant (Dependent-Samples Sign Test, S=13, p=n.s.).
@@ 5.5.2. Deployment Study - RQ2b
Similarly to the lab study, we observed little increase in the median tool satisfaction under the treatment condition (M=5.60, Q1=5.00, Q3=6.40) compared to the baseline condition (M=5.40, Q1=4.05, Q3=5.80) (Figure 7n-r). Based on median results, participants under the treatment condition were more likely to find the tool helpful (treatment: M=6.00, Q1=5.00, Q3=6.50; baseline: M=5.00, Q1=5.00, Q3=6.00), use it if it were available (treatment: M=6.00, Q1=5.00, Q3=7.00; baseline: M=5.00, Q1=4.00, Q3=6.00), and recommend the tool (treatment: M=6.00, Q1=5.00, Q3=6.50; baseline: M=5.00, Q1=3.25, Q3=6.00). They were also slightly less likely to find the tool annoying (treatment: M=2.00, Q1=2.00, Q3=4.00; baseline: M=2.50, Q1=2.00, Q3=4.00). On the other hand, they were equally likely under both conditions to find the tool useful for productivity (treatment: M=6.00, Q1=4.50, Q3=7.00; baseline: M=6.00, Q1=4.00, Q3=6.00).
@@ 5.5.3. Summary
In both studies, we observed little difference in satisfaction with Papers-to-Posts versus the baseline tool. However, participants appeared to find Papers-to-Posts more helpful than the baseline across the two studies.
@@ 5.6. Cognitive Load
To measure cognitive load, we used the NASA TLX Index questionnaire (Hart and Staveland, 1988) and summed the responses for the mental demand, temporal demand, effort, and frustration questions (Figure 8(b)a-b,d-e)=. [^30] See 5.4 for discussion of the responses to the performance question
@@ 5.6.1. Lab Study - H3
We utilized the sign test to evaluate the lab study data, as it was non-parametric data and violated the Wilcoxon signed rank test’s assumption of a symmetric distribution of the paired differences about the median. We saw no significant difference between the baseline and treatment conditions (Dependent-Samples Sign Test, S=8, p=n.s.). The median participant reported slightly lower median cognitive load with Papers-to-Posts (M=13.00, Q1=11.50, Q3=16.30) than the baseline tool (M=14.50, Q1=12.00, Q3=18.30).
We saw the largest difference in median results with respect to perceived effort (treatment: M=3.50, Q1=3.00, Q3=5.00; baseline: M=5.00, Q1=3.75, Q3=5.00). The majority of participants (11) reported using less effort with Papers-to-Posts than with the baseline tool (more: 4, equal: 5). Meanwhile, Papers-to-Posts led to little change in perceived mental demand (treatment: M=5.00, Q1=3.00, Q3=5.00; baseline: M=4.00, Q1=3.75, Q3=5.25). The small plurality of participants (9) perceived more mental demand with Papers-to-Posts (less: 8, equal: 3). We saw no change in the median result for perceived temporal demand (treatment: M=3.00, Q1=2.00, Q3=4.25; baseline: M=3.00, Q1=2.00, Q3=4.25) or frustration (treatment: M=2.00, Q1=1.00, Q3=4.00; baseline: M=2.00, Q1=2.00, Q3=3.00).
@@ 5.6.2. Deployment Study - RQ2c
Once again, we found very slightly lower median cognitive load under the treatment (M=15.00, Q1=9.50, Q3=17.50) in comparison to the baseline (M=15.50, Q1=10.75, Q3=18.75) (Figure 8(b)f-g,i-j). Considering the median participant, the treatment led to little change in perceived mental demand (M=4.00, Q1=2.00, Q3=5.50) than the baseline (M=5.00, Q1=3.25, Q3=5.75) but higher perceived effort (treatment: M=5.00, Q1=3.00, Q3=6.00; baseline: M=4.50, Q1=3.00, Q3=5.00). We found no change in the median result for perceived temporal demand (treatment: M=3.00, Q1=2.00, Q3=4.00; baseline: M=3.00, Q1=2.00, Q3=4.00) or frustration (treatment: M=3.00, Q1=2.00, Q3=5.00; baseline: M=3.00, Q1=2.00, Q3=4.75).